# Multi-State TD Target for Model-Free Reinforcement Learning

### v3

@@ I Introduction
    Introduction of the MSTD framework, incorporating the multi-state replay buffer and multi-state objective functions.
@@ VI Experiments
The empirical results demonstrate that both action-generated and action-loaded MSDDPG, as well as MSSAC, exhibit superior performance compared to both conventional multi-step and original algorithms in the HalfCheetah-v4 environment. Notably, in the Walker-v4 environment, MSDDPG achieved the highest score, while MSSAC, particularly in the action-loaded setting, surpassed other algorithms.
We also observed that the benefits of using action-loaded and action-generated settings are environment-specific. Specifically, in the Walker-v4 environment, the action-loaded setting consistently outperforms the action-generated setting, whereas the opposite trend is observed in the HalfCheetah-v4 environment.
We briefly show our results in Table 1 and other detailed information can be found in Appendix B.
@@ VII Conclusion
Appendix A: Convergence proof for Multi-state TD Target
In this part, we prove that the multi-state TD target can converge to the optimal policy under some commonly used conditions. We drive the proof based on the convergence of Q-learning mentioned in [12] and [13]. The proof of lemma 1 and the main idea of the proof of lemma 2 can be found in [13].
@@
The random process $\{\Delta_{t}\}$ taking values in $\mathbb{R}^{n}$ and defined as
@@
converges to zero under the following assumptions:
@@
    $0\leq\alpha_{t}(x)\leq 1$,$\sum_{t}\alpha_{t}(x)=\infty$ and $\sum_{t}\alpha_{t}^{2}(x)<\infty$
@@
    $\left\|\mathbb{E}\left[F_{t}(x)\right]\right\|_{\infty}\leq\gamma\left\|\Delta_{t}\right\|_{\infty},\text{ with }\gamma<1$
@@
    $\mathbf{var}[F_{t}(x)]\leq C(1+\|\Delta_{t}\|^{2}_{\infty})$, for $C>0$,
where sign $\|\|_{\infty}$ indicates the sup-norm.
@@
We define a contraction operator $\mathbf{H}^{l}$, for a function $q:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, we have
$(\mathbf{H}^{l}q)(s,a)=\sum_{y\in S}P_{l}(s,y)\big{[}\sum_{i=1}^{l}\gamma^{i-1}r_{i}+\gamma^{l}\max_{b\in\mathcal{A}}q(y,b)\big{]},$ (17)
where $l$ is a positive integer, and $P_{l}(s,y)$ indicates the transition probability between two elements in the set $\mathcal{S}$. In the context of multi-state reinforcement learning, it represents the probability that the agent will transition to state $y$ after executing $l$ actions, from state $s$, each involving a state transition. The optimal Q-function $Q^{*}$ is a fixed point of the operator $\mathbf{H}^{l}$:$Q^{*}=\mathbf{H}^{l}Q^{*}$. Meanwhile, this operator is a contraction in the up-norm:
$\|\mathbf{H}^{l}_{q_{1}}-\mathbf{H}^{l}_{q_{2}}\|_{\infty}\leq\gamma^{l}\|q_{1}-q_{2}\|_{\infty}$ (18)
@@
$=$ $\max_{x,a}\Bigg{|}\sum_{y\in\mathcal{S}}P_{l}(x,y)\big{[}\sum_{i=1}^{l}\gamma^{i-1}r_{i}+\gamma^{l}\max_{b\in\mathcal{A}}q_{1}(y,b)$
$-\sum_{i=1}^{l}\gamma^{i-1}r_{i}-\gamma^{l}\max_{b\in\mathcal{A}}q_{1}(y,b)\big{]}\Bigg{|}$
$=$ $\max_{x,a}\gamma^{l}\Bigg{|}\sum_{y\in\mathcal{S}}P_{l}(x,y)\big{[}\max_{b\in\mathcal{A}}q_{1}(y,b)-\max_{b\in\mathcal{A}}q_{2}(y,b)\big{]}\Bigg{|}$
$\leq$ $\max_{x,a}\gamma^{l}\sum_{y\in\mathcal{S}}P_{l}(x,y)\Bigg{|}\max_{b\in\mathcal{A}}q_{1}(y,b)-\max_{b\in\mathcal{A}}q_{2}(y,b)\Bigg{|}$
$\leq$ $\max_{x,a}\gamma^{l}\sum_{y\in\mathcal{S}}P_{l}(x,y)\max_{z,b}|q_{1}(z,b)-q_{2}(z,b)|$
$=$ $\max_{x,a}\gamma^{l}\sum_{y\in\mathcal{S}}P_{l}(x,y)\|q_{1}-q_{2}\|_{\infty}$
$=$ $\gamma^{l}\|q_{1}-q_{2}\|_{\infty}$ (19)
@@
For a finite MDP defined by $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, given by the multi-state TD update rule:
$Q_{t+1}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)\leftarrow Q_{t}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)+\alpha_{t}TD^{error},$ (20)
where $TD^{error}$ is:
$\frac{1}{L}\sum_{l=1}^{L}\Bigg{(}\sum_{i=1}^{l}\gamma^{i-1}\mathbf{r}_{t+i}+\max_{b\in\mathcal{A}}\gamma^{l}Q_{t}(\mathbf{s}_{t+l},\mathbf{b})-Q_{t}(\mathbf{s}_{t},\mathbf{a}_{t})\Bigg{)},$
and $Q_{t}(s_{t},a_{t})$ represents the critic output $Q_{\phi}^{\pi}(s_{t},a_{t})$ at gradient step $t$ with the input $s_{t}$, $a_{t}$ collected at time step $t$. The Q-function $Q_{t}(s,a)$ will converge to optimal Q-function $Q^{*}(s,a)$ as the following assumptions are held.
@@
    $\mathcal{S}$ and $\mathcal{A}$ are finite.
@@
    $\alpha_{t}\in[0,1]$, $\sum_{t}\alpha_{t}=\infty$, $\sum_{t}\left(\alpha_{t}\right)^{2}<\infty$.
@@
    The reward function $\mathbf{r}$ is bounded.

### v4

@@ I Introduction
    Introduction of the MSTD framework, incorporating the multi-state replay buffer and multi-state objective functions. We proved the convergence of MSTD framework, this work also shows the convergence of multi-step reinforcement learning, to our knowledge, this is the first work that proves the multi-step reinforcement learning can not converge to the optimal Q-function.
@@ VI Experiments
The experiment results and the hyperparameter for each experiment can be visited below.
[…]
hyperparameters value
hidden width 256
$\tau$ 0.005
lr 0.0004
$\gamma$ 0.99
batch size 128
$\alpha$ 0.12
TABLE 1: Hyperparameters.
hyperparameters value
hidden width 256
$\tau$ 0.005
lr 0.0003
$\gamma$ 0.99
batch size 128
$\mu$ 0
$\theta$ 0.2
$\sigma$ 0.3
TABLE 2: Hyperparameters.
@@
The empirical results demonstrate that both action-generated and action-loaded MSDDPG, as well as MSSAC, exhibit superior performance compared to both conventional multi-step and original algorithms in the HalfCheetah-v4 environment, as shown in Table 3. Specifically, in the Walker-v4 environment, the multi-step method outperforms the action-generated multi-state method but is inferior to the action-loaded multi-state method, as illustrated in Figures 3 and 4. In the HalfCheetah-v4 environment, both action-generated and action-loaded multi-state methods outperform traditional multi-step methods, with the action-generated method achieving the highest performance, as shown in Figures 5 and 6.
We also observed that the benefits of using action-loaded and action-generated settings are environment-specific. In the Walker-v4 environment, the action-loaded setting consistently outperforms the action-generated setting in terms of both convergence speed and the reward returned, as shown in Figures 3 and 4. Conversely, Figures 5 and 6 clearly show that the action-generated setting outperforms the action-loaded setting in the HalfCheetah-v4 environment.
@@ VII Conclusion
Appendix: Convergence Proof for Q-learning with MSTD
In this section, we demonstrate that Q-learning with the MSTD target converges under commonly used conditions. The convergence proof relies on two fundamental lemmas listed below. The proof of Lemma 1 and the main concept of the proof of Lemma 2 are detailed in [13] and [14], respectively.
@@
A random process
@@
converges to zero w.p.1 under the following assumptions:
@@
    The state space is finite;
@@
    $0\leq\alpha_{t}(x)\leq 1$, $\sum_{t}\alpha_{t}(x)=\infty$, and $\sum_{t}\alpha_{t}^{2}(x)<\infty$;
@@
    $\left\|\mathbb{E}\left[F_{t}(x)|P_{t}\right]\right\|_{\infty}\leq\gamma\left\|\Delta_{t}\right\|_{\infty}$ for $\gamma\in[0,1)$; and
  4. 4.
    $\operatorname{var}[F_{t}(x)|P_{t}]\leq C(1+\|\Delta_{t}\|_{\infty})^{2}$ for $C>0$,
where $P_{t}$ stands for the past at step $t$, $a_{t}$ and $F_{t}$ are allowed to depend on the past insofar as the above conditions remain valid. The notation $\|\cdot\|_{\infty}$ indicates the maximum norm and $\left\|\Delta_{t}\right\|_{\infty}=\max_{x}|\Delta_{t}(x)|$.
@@
In the context of multi-state reinforcement learning under a given policy, we consider the Q-function $q:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. We define an operator $\mathbf{H}^{l}$, for a positive integer $l$, as follows:
$(\mathbf{H}^{l}q)(s,a)=\sum_{y\in\mathcal{S}}p_{l}(y|s,a)\left[\bar{r}_{l}+\gamma^{l}\max_{b\in\mathcal{A}}q(y,b)\right],$ (17)
where $p_{l}(y|s,a)$ represents the probability of the transition to state $y$ in $l$ steps given state-action pair $(s,a)$, and $\bar{r}_{l}$ is the accumulated reward. This operator is a contraction in the sense that
$\|\mathbf{H}^{l}{q_{1}}-\mathbf{H}^{l}{q_{2}}\|_{\infty}\leq\gamma^{l}\|q_{1}-q_{2}\|_{\infty}$ (18)
for any two Q-functions $q_{1}$ and $q_{2}$.
@@
$=$ $\max_{s\in\mathcal{S},a\in\mathcal{A}}\Bigg{|}\sum_{y\in\mathcal{S}}p_{l}(y|s,a)\big{[}\bar{r}_{l}+\gamma^{l}\max_{b\in\mathcal{A}}q_{2}(y,b)-\bar{r}_{l}$
$-\gamma^{l}\max_{b\in\mathcal{A}}q_{1}(y,b)\big{]}\Bigg{|}$
$=$ $\max_{s\in\mathcal{S},a\in\mathcal{A}}\gamma^{l}\Bigg{|}\sum_{y\in\mathcal{S}}p_{l}(y|s,a)\big{[}\max_{b\in\mathcal{A}}q_{1}(y,b)-\max_{b\in\mathcal{A}}q_{2}(y,b)\big{]}\Bigg{|}.$
Additionally, we have
$\gamma^{l}\|q_{1}-q_{2}\|_{\infty}$
$=$ $\max_{s\in\mathcal{S},a\in\mathcal{A}}\gamma^{l}\sum_{y\in\mathcal{S}}p_{l}(y|s,a)\|q_{1}-q_{2}\|_{\infty}$
$=$ $\max_{s\in\mathcal{S},a\in\mathcal{A}}\gamma^{l}\sum_{y\in\mathcal{S}}p_{l}(y|s,a)\max_{z\in\mathcal{S},b\in\mathcal{A}}|q_{1}(z,b)-q_{2}(z,b)|$
$\geq$ $\max_{s\in\mathcal{S},a\in\mathcal{A}}\gamma^{l}\sum_{y\in\mathcal{S}}p_{l}(y|s,a)\Bigg{|}\max_{b\in\mathcal{A}}q_{1}(y,b)-\max_{b\in\mathcal{A}}q_{2}(y,b)\Bigg{|}.$
Comparing these two expressions implies (18). ∎
In practice, a finite number of training steps and a finite-sized replay buffer are typically used. This results in the sets $\mathcal{S}$ and $\mathcal{A}$ being finite, making $(q,\|\cdot\|_{\infty})$ a complete metric space. Furthermore, Lemma 2 demonstrates that the operator $\mathbf{H}^{l}$ is a contraction mapping. Consequently, by Banach’s Fixed Point Theorem, the operator $\mathbf{H}^{l}$ has a fixed point. Suppose there exists $Q^{f}$ such that
$(\mathbf{H}^{l}Q^{f})(s,a)=Q^{f}(s,a).$ (19)
Let $Q_{t}(s,a)$ represent the critic output $Q_{\phi}^{\pi}(s,a)$ with the inputs $s$ and $a$ at gradient step $t$. The action $a$ is either action-loaded or action-generated in the algorithms, which is related to the actor. We separate the actor’s behavior from the convergence analysis of the Q-network. Hence, the TD-error corresponding to the MSTD target (IV) becomes
$\operatorname{TD}^{\operatorname{error}}=$ $\frac{1}{L}\sum_{l=1}^{L}\Bigg{(}\sum_{i=1}^{l}\gamma^{i-1}\mathbf{r}_{t+i}+\max_{b\in\mathcal{A}}\gamma^{l}Q_{t}(\mathbf{s}_{t+l},b)$
$-Q_{t}(\mathbf{s}_{t},\mathbf{a}_{t})\Bigg{)}.$ (20)
For such a Q-learning process with the MSTD error (20), the operator $\mathbf{H}^{l}$ is specifically defined as
$(\mathbf{H}^{l}Q_{t})(\mathbf{s}_{t},\mathbf{a}_{t})$ $=\sum_{y\in\mathcal{S}}p_{l}(y|\mathbf{s}_{t},\mathbf{a}_{t})\bigg{[}\bar{\mathbf{r}}_{t,l}+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(y,b)\bigg{]}$ (21)
where $\bar{\mathbf{r}}_{t,l}=\sum_{i=1}^{l}\gamma^{i-1}\mathbf{r}_{t+i}$ for simpler notation. The convergence of Q-learning with the MSTD error is stated in the following theorem.
@@
For a finite MDP defined by $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$ where the Q-function $Q_{t}(s,a)$ is updated by the rule:
$Q_{t+1}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)=Q_{t}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)+\alpha_{t}\operatorname{TD}^{\operatorname{error}},$ (22)
where $\operatorname{TD}^{\operatorname{error}}$ is defined in (20). Then, $Q_{t}(s,a)$ converges to $Q^{f}(s,a)$ w.p.1 under the following assumptions:
@@
    $\mathcal{S}$ and $\mathcal{A}$ are finite;
@@
    $\alpha_{t}\in[0,1]$, $\sum_{t}\alpha_{t}=\infty$, $\sum_{t}\alpha_{t}^{2}<\infty$;
@@
    $\mathbf{r}_{t}$ is bounded; and